# Boosting Adversarial Transferability with Low-Cost Optimization via Maximin Expected Flatness

### v2

@@ V-B Neighborhood Conditional Sampling
The NCS begins by uniformly sampling $N$ points $\{\mathbf{x}_{0,i}\}_{i=1}^{N}$ within $B_{\xi}(\mathbf{x})$. At each iteration $t$, gradients $\nabla_{\mathbf{x}}J(\mathbf{x}_{t,i},y;\mathcal{F})$ are computed for all points, followed by a sign-based update:
$\mathbf{x}_{t+1,i}=\mathbf{x}_{t,i}-\xi\cdot\mathrm{sign}\left(\nabla_{\mathbf{x}}J(\mathbf{x}_{t,i},y;\mathcal{F})\right),$
which prioritizes directional consistency over gradient magnitudes. Updated points are then projected back to $B_{\xi}(\mathbf{x}_{0,i})$ and $B_{\gamma}(\mathbf{x})$ to enforce spatial constraints. After $T$ iterations, the final set $\{\mathbf{x}_{T,i}\}_{i=1}^{N}$ captures regions with minimal expected loss. This approach (summarized in Algorithm 1) eliminates sampling overhead by leveraging gradients, reducing computation costs while maintaining efficiency.
@@ V-C Gradient Balancing Optimization
The nested maximin structure in (4) introduces an imbalanced optimization challenge: the inner minimization requires fine-grained gradient descent across $N$ sampling points, while the outer maximization must update a single adversarial example $\mathbf{x}^{\rm adv}$ to elevate losses globally. This asymmetry arises because precise inner updates demand multiple iterations, whereas the outer optimization struggles to coordinate perturbations that simultaneously affect all minimized sub-regions. Experiments with gradient descent (1–4 inner steps) on 100 ImageNet images (Res-50→Inc-v3) validate the issue: stricter inner optimization drastically increases required convergence steps while reducing transfer attack success rates (Figure 3), confirming the inherent imbalance.
@@
$\nabla_{\mathbf{x}_{t,i}^{\rm{}^{\prime}}}J(\mathbf{x}_{t,i}^{\rm{}^{\prime}},y;\mathcal{F})\approx g_{t-1,i},$
@@
$\{g^{inner}_{t,i}\}_{i=1}^{N}=\frac{\{g_{t,i}\}_{i=1}^{N}}{\lvert\lvert\{g_{t,i}\}_{i=1}^{N}\rvert\rvert}_{1}-\mu_{inner}\cdot\{g^{inner}_{t,i}\}_{i=1}^{N},$
@@
1:Surrogate model $\mathcal{F}$ and the loss function $J$; A raw example x with ground-truth label $y$; The perturbation magnitude $\epsilon$; the maximum iterations $T$; the outer/inner decay factor $\mu_{outer}$/$\mu_{inner}$; the number of randomly sampled examples, $N$; neighborhood radius $\xi$, exploration radius $\gamma$.
@@
6:  ${\{\mathbf{x}_{t,i}\}}_{i=1}^{N}\in U_{\gamma}(\mathbf{x}^{adv}_{t})$
@@
15:  $\mathbf{x}^{adv}_{t+1}=\prod_{{B}_{\epsilon}(\mathbf{x})}[\mathbf{x}^{adv}_{t}+\alpha\cdot sign(g_{t}^{outer})]$;
@@ V-D Algorithm
The proposed framework integrates the objective of Maximin Expected Flatness, Neighborhood Conditional Sampling, and Gradient Balancing Optimization through a unified max-min bi-level algorithm summarized in Algorithm 2. The outer loop maximizes the expected flatness over adversarially identified critical sub-regions, while the inner loop approximates worst-case neighborhoods via gradient-guided sampling with momentum-stabilized direction reuse. By strategically sharing normalized gradients between inner minimization and outer maximization, the algorithm achieves an optimal balance between computational efficiency and attack effectiveness.
[…]
@@
Res-101 Inc-v3 Inc-v4 IncRes-v2 VGG-19 Dense-121 Xcept Res-50 Inc-v3 Inc-v4 IncRes-v2 VGG-19 Dense-121 Xcept
MI [21] 93.3 50.5 46.0 31.8 79.1 86.1 53.3 94.4 50.6 43.9 33.7 73.4 82.2 51.8
NI [22] 96.7 56.5 49.1 35.9 84.0 88.5 56.4 97.8 55.2 50.9 37.8 79.0 85.8 55.0
PI [4] 98.0 60.2 54.4 38.1 88.1 92.8 60.7 98.5 62.4 55.9 42.2 83.4 90.7 60.2
TPA [2] 96.2 60.5 55.0 40.4 85.9 90.5 60.7 97.2 58.1 54.2 43.2 81.4 88.8 60.9
GNP [5] 98.7 69.8 64.7 50.6 91.7 95.7 68.2 99.1 70.3 64.5 52.0 87.9 93.9 68.5
VMI [1] 98.2 72.8 68.6 57.4 91.3 95.7 73.2 98.6 73.2 68.4 57.6 88.5 93.6 70.5
VNI [1] 98.8 77.2 73.4 61.6 93.7 96.8 75.7 99.4 77.4 73.1 62.8 91.3 95.4 75.0
EMI [4] 99.3 75.0 71.2 54.3 94.8 97.2 74.3 99.8 76.8 72.6 58.5 92.9 97.2 74.0
RAP [7] 97.8 70.2 65.9 50.6 91.3 94.3 70.9 98.7 70.1 65.1 52.1 89.3 93.3 70.2
APP [6] 99.3 83.1 79.6 73.1 95.3 97.8 83.3 99.4 83.2 80.9 72.8 93.4 96.6 82.0
FEM [3] 99.8 89.4 86.3 78.8 97.4 98.7 87.7 99.8 89.1 86.4 77.6 95.4 98.0 87.3
PGN [8] 99.2 86.3 83.9 79.0 95.6 97.8 86.9 99.5 86.8 84.5 80.0 94.5 97.1 88.1
MEFH 99.7 90.6 88.7 82.7 98.3 99.4 90.4 99.7 91.2 88.2 82.2 97.0 98.4 90.0
MEFF 99.9 95.1 94.2 91.3 99.2 99.8 94.7 99.9 95.8 95.2 91.8 98.9 99.4 95.7
@@
Res-50 Res-101 Inc-v4 IncRes-v2 VGG-19 Dense-121 Xcept Res-50 Res-101 Inc-v3 IncRes-v2 VGG-19 Dense-121 Xcept
MI [21] 54.6 48.8 51.2 43.3 56.0 55.7 56.1 56.1 51.3 58.2 43.3 61.7 58.3 56.8
NI [22] 63.8 58.0 61.3 53.3 63.7 64.3 62.5 62.3 56.7 65.5 49.7 68.9 64.2 64.2
PI [4] 66.1 60.7 64.3 56.8 65.6 67.7 64.9 64.3 59.9 69.8 54.0 71.9 68.3 66.6
TPA [2] 59.4 54.2 60.4 54.2 64.3 60.9 64.7 61.6 53.3 66.7 56.1 70.7 64.2 72.0
GNP [5] 60.5 55.1 65.9 56.0 64.7 61.5 66.3 58.3 51.9 64.6 56.6 66.4 60.0 71.2
VMI [1] 69.6 65.4 71.3 64.7 70.9 72.0 71.8 70.7 66.6 77.6 66.8 76.3 74.6 73.0
VNI [1] 76.1 71.2 76.9 71.3 75.7 77.7 76.4 75.3 71.4 82.1 71.0 80.8 77.9 77.4
EMI [4] 78.8 74.3 78.5 72.2 79.5 80.5 78.6 79.0 75.3 86.5 71.9 85.4 82.1 82.7
RAP [7] 81.3 74.2 75.4 67.7 80.7 79.5 78.1 81.9 76.9 82.1 65.5 87.9 82.4 82.3
APP [6] 73.5 68.0 79.2 73.1 74.7 75.7 78.5 69.0 64.7 77.1 73.5 74.7 72.5 80.7
FEM [3] 82.1 77.1 87.2 82.1 81.8 82.1 84.5 78.3 73.9 86.0 78.9 81.6 80.4 86.9
PGN [8] 84.6 79.9 88.4 84.9 84.6 86.9 89.1 84.9 82.3 91.1 85.8 89.0 87.6 90.8
MEFH 85.8 80.8 89.4 86.1 85.2 87.4 89.9 86.2 82.4 92.8 86.1 90.1 88.5 90.8
MEFF 92.6 91.5 96.5 94.3 92.8 93.4 95.8 91.4 90.3 94.5 91.2 92.2 91.9 95.1
VI Experiments
@@ VI-A Experimental Setup
Models. We select Inception-v3 (Inc-v3) [48], Inception-v4 (Inc-v4) [57], ResNet-50 (Res-50) [9] and ResNet-101 (Res-101) [9] with TorchVision-pretrained weights [58] for generating adversarial examples. Transferability evaluation covers: (1) Standard CNNs: Inc-v3 [48], Inc-v4 [57], Res-50 [9], Res-101 [9], Inception-ResNetv2 (IncRes-v2) [57], DenseNet-121 (Dense-121) [59], VGG-19bn (VGG-19) [60] and Xception (xcept) [61]; (2) Cross-architectures: MobileNet-v2 (MobileNet) [62], PNASNet-5-Large (PNASNet-L) [63], ViT-Base/16 (ViT-B/16) [64], ViT-Large/32 (ViT-L/32) [64], PiT-S [65], MLP-Mixer [66] and ResMLP [67]; (3) Defended models: Adversarially trained Inc-v3adv [37], Inc-v3ens3 [38], Inc-v3ens4 [38] and IncRes-v2ens [38]; adversarial purification defenses HGD [41] and NRP [43]; certifiable defense RS [39].
@@
Compared Methods. We evaluate against 24 transfer-based adversarial attacks spanning three categories: (1) 5 data-driven methods covering input augmentation (DI [19], TI [20], SI [22], Admix [25], SSA [26]); (2) 4 model-driven methods comprising model ensemble (SVRE [31], CWA [32]) and surrogate refinement (SGM [30], LinBP [23]) strategies; (3) 15 optimization-driven methods including feature-based (ILA [35], NAA [24]), generative (CDA [36]), gradient stabilization (MI [21], NI [22], PI/EMI [4], VMI/VNI [1]), and flatness-enhanced (RAP [7], PGN [8], GNP [5], FEM [3], APP [6], TPA [2]) approaches. PGN [8] constitutes the state-of-the-art in optimization-driven methods.
@@
Hyper-parameters. All baseline methods use their originally reported optimal configurations unless specified. We unify common parameters across attacks: maximum perturbation $\epsilon=16/255$, iteration $T=10$, and step size $\alpha=\epsilon/T$. Sampling-based attacks employ $N=20$ samples for enhanced gradient estimation, double the typical setting in prior works. Specialized parameters retain original setups: RAP [7] uses $\alpha=1.6/255$ with $K=400$ iterations ($K_{LS}=100$ for late-start phase). Our MEF sets neighborhood radius $\gamma=2\times\epsilon$, exploration radius $\xi=0.15\times\epsilon$, outer/inner momentum coefficient $\mu_{outer}/\mu_{inner}=0.5/0.9$. Our MEF requires only one backpropagation per sample versus PGN (current SOTA)’s two. For comprehensive evaluation, we design two varients: MEFH (10 iterations, half PGN’s cost) and MEFF (20 iterations, fully matching PGN’s cost).
Evaluation Framework. In our experiment, adversarial examples generated on white-box surrogates are directly transferred to black-box targets for transfer attack success rate evaluation, with systematic analysis across three dimensions: (1) target model attributes spanning standard CNNs (Sec. VI-B), defended models (Sec. VI-C), and cross-architecture variants (Supp. Sec. -A), demonstrating attack generalization across model architectures and defense mechanisms; (2) methodological categories comparing ensemble (Supp. Sec. -B), data-driven augmentation (Sec. VI-D), and feature/generative/surrogate attacks (Sec. VI-E), highlighting superiority over existing attack paradigms; (3) method-intrinsic analysis evaluating parametric sensitivity ($\gamma/\xi$, $\mu_{outer}/\mu_{inner}$, $N/T$ in Sec. VI-F), and loss landscape via flatness measurements versus current flatness-enhanced methods (Supp. Sec. -C), validating parameter robustness and flatness gains in loss landscape.

### v3

@@ V-B Neighborhood Conditional Sampling
The NCS uniformly samples $N$ points $\{\mathbf{x}_{0,i}\}_{i=1}^{N}$ within $B_{\xi}(\mathbf{x})$. At each iteration $t$, gradients $\nabla_{\mathbf{x}}J(\mathbf{x}_{t,i},y;\mathcal{F})$ are computed for all points, followed by a sign-based update:
$\mathbf{x}_{t+1,i}=\mathbf{x}_{t,i}-\xi\cdot\mathrm{sign}\left(\nabla_{\mathbf{x}}J(\mathbf{x}_{t,i},y;\mathcal{F})\right),$ (12)
which prioritizes directional consistency over gradient magnitudes. Updated points are then projected back to $B_{\xi}(\mathbf{x}_{0,i})$ and $B_{\gamma}(\mathbf{x})$ to enforce spatial constraints. After $T$ iterations, the final set $\{\mathbf{x}_{T,i}\}_{i=1}^{N}$ captures regions with minimal expected loss. This approach eliminates sampling overhead by leveraging gradients, reducing computation while maintaining efficiency.
@@ V-C Gradient Balancing Optimization
The nested maximin structure in (11) presents an optimization imbalance: the inner minimization requires fine-grained gradient descent across $N$ sampling points, while the outer maximization must update a single adversarial example $\mathbf{x}^{\rm adv}$ to elevate losses globally. This asymmetry arises because precise inner updates demand multiple iterations, whereas the outer optimization struggles to coordinate perturbations that simultaneously affect all minimized sub-regions. Experiments with gradient descent (1–4 inner steps) on 100 ImageNet images (Res-50→Inc-v3) confirm the issue: stricter inner optimization increases required convergence steps while reducing transfer attack success rates (Figure 4).
@@
$\nabla_{\mathbf{x}_{t,i}^{\rm{}^{\prime}}}J(\mathbf{x}_{t,i}^{\rm{}^{\prime}},y;\mathcal{F})\approx g_{t-1,i},$ (13)
@@
$\{g^{inner}_{t,i}\}_{i=1}^{N}=\frac{\{g_{t,i}\}_{i=1}^{N}}{\lvert\lvert\{g_{t,i}\}_{i=1}^{N}\rvert\rvert}_{1}-\mu_{inner}\cdot\{g^{inner}_{t,i}\}_{i=1}^{N},$ (14)
@@
1:Surrogate model $\mathcal{F}$ and the loss function $J$; Raw example x with ground-truth label $y$; The perturbation magnitude $\epsilon$; the maximum iterations $T$; the outer/inner decay factor $\mu_{outer}$$\mu_{inner}$; the number of sampled examples, NN; neighborhood radius $\xi$, exploration radius $\gamma$ and $U_{\gamma}(\cdot)$ is the uniform distribution over the $\ell_{\infty}$-ball $B_{\gamma}(\cdot)$.
@@
6:  ${\{\mathbf{x}_{t,i}\}}_{i=1}^{N}\sim U_{\gamma}(\mathbf{x}^{adv}_{t})$
@@
15:  $\mathbf{x}^{adv}_{t+1}=\Pi_{{B}_{\epsilon}(\mathbf{x})}[\mathbf{x}^{adv}_{t}+\alpha\cdot sign(g_{t}^{outer})]$;
@@ V-D Algorithm
The proposed framework integrates the objective of Maximin Expected Flatness, Neighborhood Conditional Sampling, and Gradient Balancing Optimization through a unified max-min bi-level algorithm summarized in Algorithm 1. To clarify the interaction between these components, the overall workflow is visualized in Fig. 3. By strategically sharing normalized gradients between inner minimization and outer maximization, the algorithm achieves an optimal balance between computational efficiency and attack effectiveness.
VI Experiments
TABLE II: Transfer attack success rate ($\%$) comparison of gradient-stabilized attacks on normally trained models. The best results are bold and the second best are underlined.
@@ VI Experiments
Res-101 Inc-v3 Inc-v4 IncRes-v2 VGG-19 Dense-121 Xcept AVG Res-50 Inc-v3 Inc-v4 IncRes-v2 VGG-19 Dense-121 Xcept AVG
EMI [4] 99.3 75.0 71.2 54.3 94.8 97.2 74.3 80.9 99.8 76.8 72.6 58.5 92.9 97.2 74.0 81.7
APP [6] 99.3 83.1 79.6 73.1 95.3 97.8 83.3 87.4 99.4 83.2 80.9 72.8 93.4 96.6 82.0 86.9
FEM [3] 99.8 89.4 86.3 78.8 97.4 98.7 87.7 91.2 99.8 89.1 86.4 77.6 95.4 98.0 87.3 90.5
PGN [8] 99.2 86.3 83.9 79.0 95.6 97.8 86.9 89.8 99.5 86.8 84.5 80.0 94.5 97.1 88.1 90.1
ANDA [44] 99.4 81.5 80.1 74.1 90.7 94.7 82.2 86.1 99.5 82.6 81.2 76.4 91.0 92.8 83.7 86.7
FGSRA [45] 98.4 85.3 83.2 78.0 94.6 97.3 86.0 89.0 98.6 86.0 83.9 79.4 93.9 96.3 87.3 89.3
GI [46] 99.5 82.4 80.5 75.1 91.6 94.8 83.1 86.7 99.4 83.2 81.4 76.7 91.2 93.5 84.4 87.1
MUMODIG [47] 99.3 84.4 82.5 77.1 93.6 96.8 85.1 88.4 99.4 85.2 83.4 78.7 93.2 95.5 86.4 88.8
GAA [48] 98.5 82.5 81.1 75.1 91.7 95.7 83.2 86.8 98.9 83.6 82.2 77.4 92.0 93.8 84.7 87.5
FoolMix [49] 99.1 80.4 78.5 73.1 89.6 92.8 81.1 84.9 99.4 81.2 79.4 74.7 89.2 91.5 82.4 85.4
MEFH 99.7 90.6 88.7 82.7 98.3 99.4 90.4 92.8 99.7 91.2 88.2 82.2 97.0 98.4 90.0 92.4
MEFF 99.9 95.1 94.2 91.3 99.2 99.8 94.7 96.3 99.9 95.8 95.2 91.8 98.9 99.4 95.7 96.7
@@
Res-50 Res-101 Inc-v4 IncRes-v2 VGG-19 Dense-121 Xcept AVG Res-50 Res-101 Inc-v3 IncRes-v2 VGG-19 Dense-121 Xcept AVG
EMI [4] 78.8 74.3 78.5 72.2 79.5 80.5 78.6 77.5 79.0 75.3 86.5 71.9 85.4 82.1 82.7 80.4
APP [6] 73.5 68.0 79.2 73.1 74.7 75.7 78.5 74.7 69.0 64.7 77.1 73.5 74.7 72.5 80.7 73.2
FEM [3] 82.1 77.1 87.2 82.1 81.8 82.1 84.5 82.4 78.3 73.9 86.0 78.9 81.6 80.4 86.9 80.9
PGN [8] 84.6 79.9 88.4 84.9 84.6 86.9 89.1 85.5 84.9 82.3 91.1 85.8 89.0 87.6 90.8 87.4
ANDA [44] 81.1 76.6 85.2 80.6 80.8 81.9 85.2 81.6 80.9 77.7 86.3 81.0 84.2 83.8 87.5 83.1
FGSRA [45] 84.0 79.3 87.8 84.1 83.9 85.9 88.4 84.8 84.1 81.4 90.1 84.8 88.0 86.9 90.2 86.5
GI [46] 81.3 76.7 85.3 81.3 81.2 82.9 85.7 82.1 81.4 78.5 87.2 81.9 85.1 84.2 87.6 83.7
MUMODIG [47] 83.3 78.7 87.3 83.3 83.2 84.9 87.7 84.1 83.4 80.5 89.2 83.9 87.1 86.2 89.6 85.7
GAA [48] 82.1 77.6 86.2 81.6 81.8 82.9 86.2 82.6 81.9 78.7 87.3 82.0 85.2 84.8 88.5 84.1
FoolMix [49] 79.3 74.7 83.3 79.3 79.2 80.9 83.7 80.1 79.4 76.5 85.2 79.9 83.1 82.2 85.6 81.7
MEFH 85.8 80.8 89.4 86.1 85.2 87.4 89.9 86.4 86.2 82.4 92.8 86.1 90.1 88.5 90.8 88.1
MEFF 92.6 91.5 96.5 94.3 92.8 93.4 95.8 93.8 91.4 90.3 94.5 91.2 92.2 91.9 95.1 92.4
@@ VI-A Experimental Setup
Models. We select Inception-v3 (Inc-v3) [37], Inception-v4 (Inc-v4) [53], ResNet-50 (Res-50) [9] and ResNet-101 (Res-101) [9] with TorchVision-pretrained weights [54] for generating adversarial examples. Transferability evaluation covers: (1) Standard CNNs: Inc-v3 [37], Inc-v4 [53], Res-50 [9], Res-101 [9], Inception-ResNetv2 (IncRes-v2) [53], DenseNet-121 (Dense-121) [55], VGG-19bn (VGG-19) [56] and Xception (xcept) [57]; (2) Cross-architectures: MobileNet-v2 (MobileNet) [58], PNASNet-5-Large (PNASNet-L) [59], ViT-Base/16 (ViT-B/16) [60], ViT-Large/32 (ViT-L/32) [60], PiT-S [61], MLP-Mixer [62] and ResMLP [63]; (3) Defended models: Adversarially trained Inc-v3adv [29], Inc-v3ens3 [30], Inc-v3ens4 [30] and IncRes-v2ens [30]; adversarial purification defenses HGD [32] and NRP [33]; certifiable defense RS [31]; diffusion-based purifier DiffPure [34] and DensePure [35] and state-of-the-art adversarially robust models from RobustBench [64]: MIMIR [65], ARES [66] and MeanSparse [67]. (4) Multimodal Architectures (Vision-Language Models): To investigate adversarial transferability towards emerging multimodal paradigms, we evaluate VLMs including LLaVA [68], Unidiffuser [69], MiniGPT-4 [70] and Qwen2.5-VL [71].
@@
Compared Methods. We evaluate against 43 transfer-based adversarial attacks spanning three categories: (1) 11 data-driven methods covering input augmentation (DI [14], TI [15], SI [17], Admix [18], SSA [19], STM [72], USMM [73], DeCoWA [74], L2T [20], BSR [75] and OPS [21]); (2) 8 model-driven methods comprising model ensemble (SVRE [24], CWA [25]) and surrogate refinement (SGM [22], LinBP [23], AWT [76], FAUG [77], ANA [78], LL2S [79]) strategies; (3) 24 optimization-driven methods including feature-based (ILA [26], NAA [27], BFA [80], P2FA [81]), generative (CDA [28], DiffAttack [82]), gradient stabilization (MI [16], NI [17], PI/EMI [4], VMI/VNI [1], ANDA [44], FGSRA [45], GI [46], MUMODIG [47], GAA [48], FoolMix [49] ), and flatness-enhanced (RAP [7], PGN [8], GNP [5], FEM [3], APP [6], TPA [2]) approaches. PGN [8] constitutes the state-of-the-art in optimization-driven methods. We further benchmark against 4 classical query-based attacks (SimBA [13], Square [83], Boundary [12] and RayS [84]) to provide a holistic view under different threat models.
@@
Hyper-parameters. All baseline methods use their originally reported configurations unless specified (see Supp. -C). We unify common parameters across attacks: maximum perturbation $\epsilon=16/255$, and step size $\alpha=\epsilon/T$. Our MEF sets neighborhood radius $\gamma=2\times\epsilon$, exploration radius $\xi=0.15\times\epsilon$, outer/inner momentum coefficient $\mu_{outer}/\mu_{inner}=0.5/0.9$. MEF requires only one backpropagation per sample versus PGN (current SOTA)’s two. For comprehensive evaluation, we design two variants: MEFH (10 iterations, half PGN’s cost) and MEFF (20 iterations, fully matching PGN’s cost).
VI-B Effectiveness Benchmarking against Existing Attacks
VI-B1 Comparison with Gradient Stabilization Attacks
We benchmark MEF against 18 representative gradient stabilization attacks on normally trained models. We generate adversarial examples on four source models and evaluate their transferability across eight diverse target models.
Table II demonstrates the significant superiority of our framework. To maintain clarity and focus on the state-of-the-art, Table II reports methods with an average success rate exceeding 80%. Results for other classical baselines (e.g., MI [16], RAP [7]) are provided in Supp. -D (Table A-1). We compute the overall average success rate by averaging the AVG columns across all four surrogate settings. Although FEM [3] shows competitive performance on ResNet-based surrogates, PGN remains the second-best method overall with a higher global average (88.20% vs. 88.13% for FEM [3]). Notably, MEFH attains 89.93% average success, surpassing the state-of-the-art PGN [8] (88.20%) despite requiring only half the budget. Matching PGN [8]’s cost, MEFF elevates the average success rate to 94.84%, establishing MEF as the new state-of-the-art for adversarial transferability.
TABLE III: Transfer attack success rate ($\%$) comparison on diverse architectures. best results are bold and second best are underlined.
Attack Res-50 $\Longrightarrow$ Inc-v3 $\Longrightarrow$
MobileNet PNASNet-L ViT-B/16 ViT-L/32 PiT-S MLP-Mixer ResMLP MobileNet PNASNet-L ViT-B/16 ViT-L/32 PiT-S MLP-Mixer ResMLP
FEM [3] 94.0 87.0 43.1 35.0 58.1 60.2 66.0 81.1 78.2 41.3 30.4 53.2 59.7 58.6
PGN [8] 91.6 88.6 51.4 45.8 58.7 66.7 70.2 84.9 82.6 50.1 40.6 60.2 66.1 65.6
FGSRA [45] 88.8 87.0 49.8 45.6 57.1 64.5 70.9 83.7 81.7 48.1 39.6 59.6 64.2 65.2
MUMODIG [47] 89.5 86.3 48.8 44.1 56.9 66.3 69.0 81.3 81.1 48.6 39.0 59.5 65.7 65.1
MEFH 95.2 90.3 51.9 46.0 63.3 66.9 70.5 85.2 83.9 51.7 42.2 60.4 67.3 66.3
MEFF 98.0 95.2 56.0 46.9 71.1 71.7 77.4 92.7 90.8 56.2 45.3 71.4 69.4 73.3
VI-B2 Evaluation on Diverse Network Architectures
To assess cross-architecture transferability, we evaluate attack success rates on seven heterogeneous models, including MobileNet [58], PNASNet [59], Vision Transformers (ViT) [60, 61], and MLP-Mixers [62, 63]. Table III compares MEF against the top-4 performing baselines (full comparison with 12 baselines is in Supp. -G). Despite the significant inductive bias gap, MEFF consistently outperforms the state-of-the-art PGN [8], achieving a remarkable 46.9% success rate on ViT-L/32 (vs. PGN’s 45.8%) and 77.4% on ResMLP (vs. PGN’s 70.2%). This confirms that optimizing zeroth-order flatness effectively extracts geometry-invariant features that transfer across fundamentally different architectural paradigms.
TABLE IV: Transfer attack success rate (%) comparison on ensemble surrogates. Best results are bold and second best are underlined.
Attack Res-50 + Inc-v3 $\Longrightarrow$
Res-101 Inc-v4 IncRes-v2 Inc-v3adv Inc-v3ens3 Inc-v3ens4 IncRes-v2ens MobileNet PNASNet-L ViT-B/16 ViT-L/32 PiT-S MLP-Mixer ResMLP
FEM [3] 98.7 92.0 88.3 68.3 65.9 62.9 46.9 92.5 89.0 59.0 41.3 74.1 66.0 77.0
PGN [8] 97.8 91.9 89.4 79.5 78.2 77.5 65.6 93.9 91.9 68.7 50.2 76.9 74.5 81.8
FGSRA [45] 94.9 90.3 87.4 78.9 76.4 75.2 66.3 92.6 90.9 66.5 49.1 76.1 72.5 81.2
MUMODIG [47] 95.5 89.6 86.0 77.0 76.0 76.9 64.5 90.1 90.2 66.8 48.4 75.8 73.9 80.9
MEFH 98.9 93.2 90.2 80.2 79.8 80.5 67.0 94.1 92.4 69.5 51.7 77.8 75.3 82.1
MEFF 99.4 96.6 95.1 85.7 83.3 82.6 70.4 96.7 95.1 75.4 55.2 83.8 80.4 89.0
TABLE V: Transfer attack success rate (%) comparison with input augmentation attacks. The best results are bold.
Attack Res-50 $\Longrightarrow$ Inc-v3 $\Longrightarrow$
IncRes-v2 Inc-v3ens3 Inc-v4ens4 IncRes-v2ens ViT-L/32 MLP-Mixer AVG IncRes-v2 Inc-v3ens3 Inc-v4ens4 IncRes-v2ens ViT-L/32 MLP-Mixer AVG
TI [15] 51.1 42.0 39.1 30.2 34.4 55.6 42.1 52.0 37.0 37.2 26.0 28.8 53.7 39.1
MEFH+TI 81.7 77.4 77.8 68.2 57.3 71.1 72.2 86.4 79.6 79.5 67.4 48.7 68.3 71.6
MEFF+TI 94.1 83.5 82.6 73.6 62.5 76.0 78.7 94.5 89.0 88.3 78.8 56.3 75.6 80.4
STM [72] 84.9 52.0 49.1 32.9 37.5 66.4 53.8 85.6 54.2 56.4 32.7 34.9 63.8 54.6
USMM [73] 84.8 51.7 48.8 32.6 36.9 66.0 53.5 85.4 53.4 56.3 32.3 34.9 63.3 54.3
DeCowA [74] 90.1 59.4 56.7 40.6 44.3 70.2 60.2 92.9 62.3 63.9 40.2 42.8 71.6 62.3
L2T [20] 89.6 56.8 53.8 37.5 41.6 69.7 58.2 90.3 58.5 61.1 37.0 40.1 69.3 59.4
BSR [75] 89.4 57.9 55.0 38.6 43.1 69.9 59.0 91.4 59.5 62.2 38.2 40.1 70.0 60.2
OPS(10,5,5) [21] 96.2 84.3 81.9 72.3 63.7 77.7 79.3 95.9 90.7 88.8 79.4 57.1 76.3 81.4
OPS(10,30,30) [21] 98.5 90.6 89.2 83.7 81.1 98.1 90.2 98.7 93.1 93.2 86.6 75.3 91.9 89.8
MEFH 83.7 64.5 63.3 50.4 40.8 65.3 61.3 86.1 65.1 65.4 45.4 35.2 65.5 60.4
MEFF 90.3 69.2 65.1 52.2 44.9 71.7 65.6 94.3 69.0 69.6 47.9 40.3 69.4 65.1